A conical tank that is 5 meters high has a radius of 2 meters, and is filled with a liquid that weights 800 kg per cubic meter. How much work is done in discharging all the liquid at point 3 meters above the top of the tank?
1
Expert's answer
2020-04-15T18:05:11-0400

Let we have an inverted cone.

Put the origin of the coordinate system at the center of the top of the tank. Have coordinates measure depth (so y = 0 at the center of the top of the tank, y = 5 at the bottom of the tank).

From similar triangles


"{x \over 5-y}={2 \over 5}=>x=2-{2 \over 5}y"

Segment the liquid into a stack of "disks", each with thickness "dy". One such "disk" at height "y" from the bottom of the tank has Volume


"dV=\pi x^2dy=\pi(2-{2 \over 5}y)^2 dy"

The mass of this "disk" is


"dm=\rho dV=800\pi(2-{2 \over 5}y)^2dy=\pi (3200-1280y+128y^2)dy"

The vertical distance this disk must travel is "(y+3)"

The weight (force due to gravity) of the disk is "dm\cdot g," where "g" is the gravitational constant 

"g=9.81\ m/s^2."


"Work=\pi\displaystyle\int_{0}^59.81\cdot(3200-1280y+128y^2)(y+3)dy="

"=1255.68\pi\big[{y^4 \over 4}-{7y^3 \over 3}-{5y^2 \over 2}+75y\big]\begin{matrix}
 5 \\
 0
\end{matrix}=44472\pi\approx139713(J)"